I'm really confused trying to solve this equation. I keep getting different answers. I think the bx is confusing me.
3bx + 9 = 4(3x-b)+2 solve for x

Found 2 solutions by usyim88hk, Scorinitron:
Answer by usyim88hk(158)

3bx+9 = 4(3x-b)+2
First, multiply 4 with (3x-b)
3bx+9 = 12x-4b+2
Now munis 9 from both sides and minus 12x from both sides
3bx-12x = -4b-7
Factor 3bx-12x
3x(b-4) = -4b-7
Divide both sides by (b-4)
3x = (-4b-7)/(b-4)
Now divide both sides by 3
x = (-4b-7)/(3b-12)

Answer by Scorinitron(19)

I'm going to be real. This was hard...and confusing but here's what I did.
The Problem
3bx+9=4(3x-b)+2
____________________
The Steps:
______________________
Multiply 4 by each term inside the parentheses.
3bx+9=12x-4b+2

Since 12x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 12x from both sides.
3bx+9-12x=-4b+2

Factor out the GCF of 3x from each term in the polynomial.
3x(b)+3x(-4)=-4b-7

Factor out the GCF of 3x from 3bx-12x.
3x(b-4)=-4b-7

Divide each term in the equation by (b-4).
(3x(b-4))/(b-4)=-(4b)/(b-4)-(7)/(b-4)
Simplify the left-hand side of the equation by canceling the common factors.

3x=-(4b)/(b-4)-(7)/(b-4)

Simplify the right-hand side of the equation by simplifying each term.
3x=(-4b-7)/(b-4)

Divide each term in the equation by 3.
(3x)/(3)=(-4b-7)/(b-4)/(3)

Simplify the left-hand side of the equation by canceling the common factors.
x=(-4b-7)/(b-4)/(3)

Simplify the right-hand side of the equation by simplifying each term.
x=(-4b-7)/(3(b-4))